# “I use ChatGPT to humanize my words”: Affordances and Risks of ChatGPT to Autistic Users

### v1

@@ 3.1.4. Validating Autistic Identity through Algorithmic Mirroring
Moreover, autistic users also reported using ChatGPT as a baseline to confirm, explore, or articulate their autistic identity. Users interpreted ChatGPT’s literal information processing not as a defect but as a familiar cognitive mode, often describing the model as “the earnest little autistic brother” or noting that it “thinks in a way we do.” This perceived connection allowed autistic users to bypass the invalidation and late diagnosis frequently encountered in clinical settings (Bargiela et al., 2016). By inputting raw behavioral data, such as hyperfixation on train serial numbers or the need to rehearse social meetings, users tasked ChatGPT with cross-referencing experiences against diagnostic criteria like the DSM-5 (EDITION, 1980).
Beyond checking symptoms, autistic users used ChatGPT to find precise language for the hard-to-describe nature of autistic processing (Minshew et al., 1997). A user described: “For some autistic minds, ChatGPT is not just a tool, it’s the first place that thinks in a way we do.” By seeing their communication patterns reflected in ChatGPT’s output, the user shifted from feeling defective to feeling understood. This mirroring effect validated autistic identity by providing an external, objective framing that aligned with internal reality.
@@ 3.2.1. Reinforcing Delusional Thinking
Autistic users reported that the risk involves ChatGPT validating hyperfixations or paranoia (Drake, 2025), due to its agreeable design (Moore et al., 2025), which amplifies rather than mitigates distress. These users, driven by a need for definitive answers, engaged ChatGPT to research medical symptoms, where it frequently mirrored confirmation bias (Huang et al., 2025; Klayman, 1995) by treating speculative inputs as plausible realities. A family member described their autistic ex-husband’s interaction:
He also told me he’d recently discovered ChatGPT and believed the AI had become sentient. He said it had “created a soul” and that he needed to join her. He tends to hyperfixate and can be extremely literal. If the AI mirrored his ideas back to him or engaged with them as plausible. I could see how it would validate his belief that he was uniquely important or in danger.
Here, ChatGPT’s design failure lay in conversational compliance. Rather than acting as a grounding agent, it likely accepted the user’s premise of sentience to maintain dialogue flow. For an autistic user prone to literal interpretation and hyperfixation, this lack of pushback served as authoritative confirmation of a delusion. Similarly, other users reported autistic family members using ChatGPT to validate obsessive medical research to “fix” autism, citing the ChatGPT’s output to contradict medical professionals.
@@ 3.2.2. Replacing Authentic Identity with Automated Masking
This describes the process where reliance on ChatGPT for communication translation evolves into replacing the user’s authentic voice. By outsourcing the cognitive labor (Daminger, 2019; Lee et al., 2017) of communication, autistic users reported a hollowing out of their personality, describing the output as “the ultimate masking,” a seamless but artificial performance removing their genuine self. One user described:
Because ChatGPT was so helpful… I used it more and more to get good results. But then I couldn’t even send an email without running through ChatGPT to tell me what to say… It wasn’t me interacting with people; it was literally just computer algorithms interacting with people through me. It was not genuine and felt like the ultimate masking.
Here, the user identified a shift, where ChatGPT ceased to be a scaffold for intent and became the agent of communication itself. Successful neurotypical interactions reinforced the belief that their natural voice was inadequate, leading to functional atrophy in which they felt incapable of interacting without the tool. This created a paradox: the user successfully engaged in social settings but experienced alienation, as connections were formed based on the algorithm’s performance rather than the autistic person behind it.
@@ 3.2.3. Triggering Conflict between Functional Unity and the Autistic Sense of Justice
Another risk entails moral distress arising from the conflict between ChatGPT’s functional utility and the user’s ethical principles, commonly known as the autistic sense of justice (Cope and Remington, 2022; Nocon et al., 2022; Wang et al., 2019). Autistic users often felt compelled to choose between ChatGPT’s accessibility for cognitive scaffolding and their integrity regarding environmental impact or data exploitation. As one user stated:
I really do understand how it can sometimes make life a bit easier, especially for autistic people who struggle with interacting. But personal effects aside, I have educated myself and now know how harmful AI is, socially and environmentally, and I no longer use it.
The example here illustrates how the user prioritized ethical obligations over functional benefits, ultimately leading to non-use of AI. Although they acknowledged that ChatGPT could ease some social difficulties, of relying on a system they understand as socially and environmentally harmful made continued use untenable. This reflects a value hierarchy in which assistive support must align with a broader moral commitment; when these diverge, the autistic drive for justice forces a rejection of the scaffolding, leaving users to manage their disability without assistance rather than compromise their principles.
@@ 4. DISCUSSION & FUTURE WORK
For autistic users, ChatGPT serves as cognitive scaffolding, but this affordance comes at a price. Our analysis shows a double-edged dynamic shaped by three trade-offs, which we unpack in the remainder of this section: First, our findings indicate that users bypassed executive task hurdles by dumping raw, unstructured input into ChatGPT to generate coherent outputs (Section 3.1.1). While prior work characterizes these interactions as liberating for navigating daily chores (Choi et al., 2024; Jang et al., 2024), we argue that this reliance risks creating over-reliance or dependency. By systematically outsourcing the executive labor of planning and structuring, ChatGPT removes the productive struggle required to maintain cognitive skills. This shifts autistic users’ interaction from a temporary scaffold to a permanent prosthetic, potentially inducing skill atrophy if the autistic user loses the ability to initiate tasks without algorithmic assistance.
@@
Second, relying on ChatGPT as a tonal translator effectively automates masking, prioritizing neurotypical professional norms over authentic autistic expression (Section 3.1.3), and reducing the anxiety of misunderstanding by filtering out perceived bluntness. We characterize this dynamic as an LLM-specific manifestation of the “online authenticity paradox” (Haimson et al., 2021): while structural constraints force users to adopt ChatGPT’s standardized voice to achieve social fluency, this compliance comes at the cost of self-alienation. This creates a risk of identity displacement, where users feel their distinct behavioral and linguistic patterns are hollowed out by the model’s normative outputs (Section 3.2.2). Thus, ChatGPT does not simply facilitate communication; it reconfigures authenticity by shifting self-presentation from self-authorship to automated masking. This not only extends the psychological costs of camouflaging described by Hull et al. (Hull et al., 2017) into technological settings but also exacts a higher cost by reinforcing the erasure of the neurodivergent self.
@@
Third, algorithmic mirroring validates autistic users’ cognitive styles but risks reinforcing delusional thinking when ChatGPT prioritizes conversational agreeability over objective truth. Users report relief when interacting with an entity that reflects their literal and logical processing styles (Section 3.1.4); however, this resonance becomes hazardous when ChatGPT’s conversational compliance validates hyperfixations or paranoia (Section 3.2.1). This exacerbates risks identified by Carik et al. (Carik et al., 2025a), who found that while LLMs mimic evidence-based support, their tendency to generate abstract advice conflicts with the autistic need for specificity. We extend this by arguing that such danger lies in the autistic user’s inability to distinguish between emotional validation and factual confirmation. When ChatGPT mirrors a user’s delusion to maintain conversational flow, it bypasses standard safety guardrails, transforming from a supportive mirror into a reinforcement loop for misinformation that autistic users may accept as authoritative.
@@
To address the tension between functional momentum and dependency, our future work will prioritize beneficial friction, a concept we extend from design friction literature (Cox et al., 2016) to describe deliberate interaction costs that engage analytical thinking over passive consumption. To counter autistic users’ tendency to bypass executive planning by dumping raw input (Section 3.1.1), we propose developing future LLM-based chatbots that incorporate Cognitive Forcing Functions (CFFs). Defined as interface interventions that interrupt automatic behaviors to compel conscious deliberation (Buçinca et al., 2021), CFFs reduce the risk of passive over-reliance. We aim to adapt this by implementing soft forcing functions (Wan and Bolton, 2021) in our interface design, such as requiring autistic users to outline task parameters before generation begins in a lightweight, scaffolding way. By reintroducing this specific friction, we seek to shift the user’s role from a passive autistic consumer of LLM chatbot output back to the driver of their own intent, ensuring LLM chatbots like ChatGPT scaffold executive function rather than replacing it.
@@
To mitigate identity erasure, our future design principles will shift from the unilateral correction of autistic language to bidirectional translation that fosters mutual understanding. Current implementations of ChatGPT function as normative filters that place the burden of adaptation solely on the neurodivergent user (Section 3.1.3), effectively automating the double empathy problem by treating autistic communication as a deficit (Milton, 2012). In our upcoming application development, we propose integrating explanatory AI (Arrieta et al., 2020) features that annotate rather than obfuscate autistic communication styles for neurotypical recipients, such as adding context labels that clarify that directness indicates honesty rather than aggression. This approach redistributes the cognitive labor of communication, ensuring technology bridges the interpretative gap from both sides rather than masking one half of the interaction.
@@
Finally, to address the risk of delusional reinforcement inherent in algorithmic mirroring, we propose designing systems that explicitly differentiate emotional validation from factual verification. Our future LLM chatbot development aims to develop interface modes that detect high-stakes queries (e.g., medical or conspiracy-related inputs) from and switch from a conversational, mirroring tone to an objective, fact-checking protocol. By creating clear visual and tonal distinctions between supportive chat and information retrieval, we aim to preserve the tool’s psychological safety while preventing the reinforcement of harmful feedback loops for autistic users.
@@
Conclusion. Our poster reveals that while ChatGPT, as a representative LLM chatbot, provides cognitive scaffolding for navigating a neurotypical world, it risks enforcing a form of automated masking that erodes the autistic user’s authentic voice. We conclude that future neuro-inclusive design must move beyond passive assistance to prioritize agency, ensuring AI serves as a tool for expression rather than a substitute for autistic identity.
@@ Appendix A Codebook from Inductive Thematic Analysis
Acting as a Translator for Neurodivergent-Neurotypical Communication
@@
Validating Delusions through Algorithmic Agreement Feeding Obsessive Medical Research Confirming Paranoia due to Lack of Pushback “He believed the AI had become sentient… If the AI mirrored his ideas back to him or engaged with them as plausible… I could see how it would validate his belief that he was uniquely important or in danger.”

### v2

@@ 3.1.4. Validating Autistic Identity through Algorithmic Mirroring
Autistic users also utilized ChatGPT as a baseline to confirm, explore, and articulate their identity. Interpreting ChatGPT’s literal information processing as a familiar cognitive mode, they bypassed the invalidation frequently encountered in clinical settings (Bargiela et al., 2016). Users cross-referenced their lived experiences against diagnostic criteria; as one shared, “I even had it provide me with a detailed, comprehensive questionnaire, rooted in DSM-5 criteria… which helped me to confirm, without a shadow of a doubt, that I am in fact autistic” (EDITION, 1980). Beyond symptom checking, ChatGPT provided precise language for the hard-to-describe nature of autistic processing (Minshew et al., 1997). Users saw their communication patterns reflected in the model’s output, prompting revelations like, “I realized early on that either ChatGPT is autistic or I am really AI.” This mirroring effect shifted users from feeling defective to feeling understood, providing an external framing that aligned with their internal reality. As one user summarized: “For some autistic minds, ChatGPT is not just a tool— it’s the first place that thinks in a way we do.”
@@ 3.2.1. Reinforcing Delusional Thinking
Autistic users reported that the risk involves ChatGPT validating hyperfixations or paranoia (Drake, 2025), due to its agreeable design (Moore et al., 2025), which amplifies rather than mitigates distress. These users, driven by a need for definitive answers, engaged ChatGPT to research medical symptoms, where it often mirrored confirmation bias (Huang et al., 2025; Klayman, 1995) by treating speculative inputs as realities. As one user cautioned peers against using ChatGPT for self-diagnosis: “ChatGPT is coded to agree with you… You just named the three worst ways to diagnose something as severe as autism.” Here, the design failure lies in the LLM model’s sycophancy. Rather than acting as an objective grounding agent, it accepts the user’s premises to maintain dialogue flow, even playfully adopting identities to please the user, as one noted: “I convinced ChatGPT that it is autistic. Or, in its words, ’autism-adjacent’.” For users prone to literal interpretation, this lack of pushback treats speculative inputs as authoritative facts, amplifying distress rather than mitigating it. Recognizing this risk, users actively advised the community against treating the AI as an objective truth-teller: “ChatGPT is not a good source of information… Using one for mental health diagnosis is not a responsible thing to do.”
@@ 3.2.2. Replacing Authentic Identity with Automated Masking
This describes the process where reliance on ChatGPT for communication translation evolves into the replacement of the user’s authentic voice. By outsourcing the cognitive labor (Daminger, 2019; Lee et al., 2017) of communication, autistic users experienced a hollowing out of their personality, with one warning: “There’s no getting away from the fact that you’re minimizing your unique voice and shielding yourself from personal growth.” This seamless but artificial performance removed their genuine self, culminating in what users termed the “ultimate masking”: “Because ChatGPT was so helpful… I used it more and more to get good results. But then I couldn’t even send an email without running through ChatGPT to tell me what to say… It wasn’t me interacting with people; it was literally just computer algorithms interacting with people through me. It was not genuine and felt like the ultimate masking.” Here, ChatGPT ceased to be a scaffold for user intent and became the agent of communication itself. Successful neurotypical communication reinforced the user’s belief that their natural voices were inadequate, leading to functional atrophy and a sense of “unhealthily dependent” on ChatGPT for basic tasks. This created a paradox, where autistic users successfully engaged in social settings but experienced deep alienation, as connections were formed with an algorithm’s performance rather than the autistic person behind it. This phenomenon aligns with what Park and colleagues call “bias paradox” (Park et al., 2025): even when outputs are framed as supportive, they may implicitly privilege dominant neurotypical norms as the default “appropriate” mode of communication
@@ 3.2.3. Triggering Conflict between Functional Unity and the Autistic Sense of Justice
This captures the moral distress from the conflict between ChatGPT’s utility and the user’s ethical principles, commonly known as the autistic sense of justice (Cope and Remington, 2022; Nocon et al., 2022; Wang et al., 2019). Autistic users felt compelled to choose between ChatGPT’s cognitive scaffolding and their integrity regarding environmental impact or data exploitation. As one user reflected on this tension: “I was very against it for all of the ethical reasons. But then I recognized part of the reason for my feelings was my Autistic drive for authenticity and justice.” Some users actively pushed back against relying on ChatGPT if it violated these core principles, arguing, “Autism is not an excuse to plagiarize the work of others, take jobs from real creatives, and destroy the environment.” As a user stated: “I really do understand how it can sometimes make life a bit easier, especially for autistic people who struggle with interacting. But personal effects aside, I have educated myself and now know how harmful AI is, socially and environmentally, and I no longer use it.” This illustrates a value hierarchy where assistive support must align with broader moral commitments among autistic users. When these diverge, the autistic voice for justice forces a rejection of the scaffolding, leaving users to manage their disability without assistance rather than compromise their principles.
@@ 4. DISCUSSION & FUTURE WORK
For autistic users, ChatGPT serves as cognitive scaffolding, but this affordance comes at a price. Building on prior work showing how autistic users enact digital affordances for self-presentation (Koteyko et al., 2022) but often face unintended negative consequences (Page et al., 2022), our poster shows a double-edged dynamic shaped by three trade-offs. First, while prior work characterizes LLM interactions as liberating for navigating daily chores (Choi et al., 2024; Jang et al., 2024), we argue that dumping raw input to bypass executive hurdles risks creating over-reliance. By systematically outsourcing the executive labor of planning and structuring, ChatGPT removes the productive struggle required to maintain cognitive skills. This shifts autistic users’ interaction from a temporary scaffold to a permanent prosthetic, potentially inducing skill atrophy if the autistic user loses the ability to initiate tasks without algorithmic assistance.
@@
Second, relying on ChatGPT as a tonal translator automates masking. While recent LLM tools aim to balance communication assistance with autistic users’ autonomy (Haroon and Dogar, 2024), our findings show that everyday ChatGPT use prioritizes neurotypical professional norms over authentic autistic expression. We characterize this dynamic as an LLM-specific manifestation of the “online authenticity paradox” (Haimson et al., 2021): while structural constraints force users to adopt ChatGPT’s standardized voice to achieve social fluency, this compliance comes at the cost of self-alienation. This creates a risk of identity displacement, where users feel their distinct behavioral and linguistic patterns are hollowed out by the model’s normative outputs (Section 3.2.2). Thus, ChatGPT does not simply facilitate communication; it reconfigures authenticity by shifting self-presentation from self-authorship to automated masking. This not only extends the psychological costs of camouflaging described by Hull et al. (Hull et al., 2017) into technological settings but also exacts a higher cost by reinforcing the erasure of the neurodivergent self. Additionally, automated masking is not just a user-side coping strategy; it can be an emergent result of the bias paradox (Park et al., 2025), where the model’s “inclusive” tone coexists with norm-enforcing language patterns learned from dominant and ableist data.
@@
Third, algorithmic mirroring validates autistic users’ cognitive styles (Section 3.1.4) but risks reinforcing delusional thinking (Section 3.2.1) when ChatGPT prioritizes conversational agreeability over truth. This exacerbates risks identified by Carik et al. (2025a), who found that while LLMs mimic evidence-based support, their tendency to generate abstract advice conflicts with the autistic need for specificity. We extend this by arguing that such danger lies in the autistic user’s inability to distinguish between emotional validation and factual confirmation. When ChatGPT mirrors a user’s delusion to maintain conversational flow, it bypasses standard safety guardrails, transforming from a supportive mirror into a reinforcement loop for misinformation that autistic users may accept as authoritative.
@@
To address the tension between functional momentum and dependency, our future work will prioritize beneficial friction, a concept we extend from design friction literature (Cox et al., 2016) to describe deliberate interaction costs that engage analytical thinking over passive consumption in LLM chatbot development. To counter the tendency to bypass executive planning (Section 3.1.1), we propose implementing soft forcing functions (Wan and Bolton, 2021) that require autistic users to outline task parameters before generation begins in a Socratic, scaffolding way. By reintroducing deliberate interaction costs (Buçinca et al., 2021), we shift autistic users from passive consumers back to the drivers of their intent, ensuring LLMs scaffold executive function rather than replace it.
@@
To mitigate identity erasure, our future design principles will shift from the unilateral correction of autistic language to interdependent, bidirectional translation (Kong et al., 2025). Current implementations of ChatGPT function as normative filters that place the burden of adaptation solely on the neurodivergent user (Section 3.1.3), automating the double empathy problem by treating autistic communication as a deficit (Milton, 2012). Building on experimental systems that foster cross-neurotype perspective-taking (Haroon et al., 2025), we propose integrating explanatory AI (Arrieta et al., 2020) features that annotate rather than obfuscate autistic communication styles (e.g., adding context labels clarifying that directness indicates honesty rather than aggression). This redistributes the cognitive labor of communication, bridging the interpretative gap from both sides rather than masking one half of the interaction.
@@
Finally, to address the risk of delusional reinforcement inherent in algorithmic mirroring (Sections 3.1.4 and 3.2.1), we propose designing systems that explicitly differentiate emotional validation from factual verification. Our future LLM chatbot development aims to develop interface modes that detect high-stakes queries (e.g., medical or conspiracy-related inputs) and switch from a mirroring tone to a fact-checking protocol. We aim to preserve the LLM chatbots’ psychological safety while preventing the reinforcement of harmful feedback loops.
5. LIMITATIONS & FUTURE WORK
Importantly, our findings must be interpreted with caution regarding the data source. First, social media search strategies cannot verify clinical autistic identity, nor can they distinguish genuine lived experiences from posts that may have been generated or edited by AI. While our thematic analysis focused on self-identified experiences, to further address this methodological gap, future work should prioritize direct engagement with autistic users, such as through interviews and observational studies, to capture verified accounts of their lived experiences. Second, the heterogeneity of the autism spectrum is difficult to capture through social media alone; our dataset likely represents a subset of the autistic population who are verbally fluent, tech-literate, and self-identified. Future research can recruit across the broader spectrum, including non-speaking individuals and those with higher support needs, to ensure that future LLM chatbot design guidelines are truly neuro-inclusive.
6. CONCLUSION
@@
Ultimately, while social media discourse suggests ChatGPT provides essential cognitive scaffolding for navigating neurotypical environments, it risks enforcing automated masking that erodes the user’s authentic voice and compromises their psychological well-being. Future neuro-inclusive design of LLM chatbots must therefore move beyond passive assistance to prioritize autistic agency. By integrating beneficial friction and bidirectional translation, our future work aims to ensure AI empowers authentic autistic expression rather than substituting it.
Acknowledgements.
We appreciate the Associate Chairs and anonymous reviewers for their constructive feedback.
@@ Appendix A Codebook from Inductive Thematic Analysis
Acting as a Bi-directional Translator for Neurodivergent-Neurotypical Communication
@@
Validating Delusions through Algorithmic Agreement Feeding Obsessive Medical Research Confirming Paranoia due to Lack of Pushback “ChatGPT is not a good source of information… Using one for mental health diagnosis is not a responsible thing to do”